# Understanding In-Context Learning with a Pelican Soup Framework

### v1

@@ Appendix E Detailed Gengeration Process of the LM Training Data in Calcutec
           Reject the sampled $X\cup\{x^{\prime}\}$. {We don’t want a premise that entails everything.}
@@
        Sample a $g\in M^{\prime}$ such that the proof tree for $X^{\prime}\models g$ contains $s$ and its depth $>d_{min}$. {We use $d_{min}=4$ in our experiments.}
@@
     $D\leftarrow$ $D$.replace($s^{\prime}$, $r_{b}$)
@@
Algorithm 1 Pseudo code for the generation process of an Calcutec document used for training.
Appendix F Perturbations in Calcutec
We apply two types of perturbations over the reasoning steps in Calcutec described in §4:
@@ Appendix F Perturbations in Calcutec
Additionally, when flatting the proof trees with topological sort, we break ties randomly. We also randomize the order of the symbols in the antecedents.
[…]
@@
Appendix G Extra Experiments with Calcutec
G.1 Additional Setups
@@
Based on Assumption 2.9 and Assumption 2.5, input-label pairs of a downstream task corresponds to prefix-reference pairs in a paragraph. To examine whether the trained model can generalize well when the induction process for the label is different from the induction process for the pronoun in the training data, we generate a training dataset where all the pronouns are induced from the premise with a left-branching proof tree with a depth equal to 2 (Figure 3(a)), while the test data contains samples whose labels are induced from the input with balanced trees (Figure 3(b)).
@@
As described in §F, we apply some random perturbations on the proving process. We also experiment with the setup where we do not apply any perturbations.
@@
As described in §E, we rewrite the first step of the proof. We also experiment with the setup where we do not rewrite the first step.
@@ G.2 Results and Discussion
Figure 8 show that deeper models have better ICL performance. It is aligned with the real-world observation that scaling helps model performance.
[…]
@@ Appendix H Hyper-parameters
(a) balanced, length = $4$, $y_{1},y_{2}$
(b) branching, length = $4$, $y_{1},y_{2}$
(c) branching, length = $3$, $y_{1},y_{2}$
(d) balanced, length = $3$, $y_{3},y_{4}$
(e) branching, length = $4$, $y_{3},y_{4}$
(f) branching, length = $3$, $y_{3},y_{4}$
Figure 5: In-context learning accuracy with Calcutec when using different verbalizers ($y_{1},y_{2}$ or $y_{3},y_{4}$) and input lengths (3 or 4). The dotted lines represent the performance of unseen combinations described in §4.2, while the different colors represent the number of formulas each class ($v_{+}$ or $v_{-}$) is associated to. The main lines represent the average accuracy of 5 tasks. We plot the performance of each task in lighter colors.
(a) balanced, length = $4$, $r_{1},r_{2}$
(b) branching, length = $4$, $r_{1},r_{2}$
(c) branching, length = $3$, $r_{1},r_{2}$
(d) balanced, length = $4$, $r_{3},r_{4}$
(e) branching, length = $4$, $r_{3},r_{4}$
(f) branching, length = $3$, $r_{3},r_{4}$
@@
(a) balanced, length = $4$, $r_{1},r_{2}$
(b) branching, length = $4$, $r_{1},r_{2}$
(c) branching, length = $3$, $r_{1},r_{2}$
(d) balanced, length = $4$, $r_{3},r_{4}$
(e) branching, length = $4$, $r_{3},r_{4}$
(f) branching, length = $3$, $r_{3},r_{4}$
@@
(a) balanced, length = $4$, 3 layers
(b) balanced, length = $4$, 4 layers
(c) balanced, length = $4$, 5 layers
(d) branching, length = $4$, 3 layers
(e) branching, length = $4$, 4 layers
(f) branching, length = $4$, 5 layers
@@
(a) 5-digit
(b) 6-digit
Figure 9: The distribution of the number of reasoning steps in the dataset when some of them are dropped at different probability. Each number is the average over 5 datasets generated with different random seeds.
Appendix I Formal Description of the Digit Addition Data
[…]
(a) 5 digits
(b) 6 digits
Figure 10: The accuracy of the models for the addition tasks. The x-axis represents the probability at which we drop each reasoning step in the training data independently. The solid line represents the ratio of testing samples where the model can output the exact answer, while the dashed line represents the character-level accuracy. The results are the average of 5 random seeds.
Figure 11: The exact accuracy (y-axis, solid points) and digit-level accuracy (y-axis, hollow points) versus validation loss (x-axis) for a 5-digit addition task.
Figure 12: The exact accuracy (y-axis, solid points) and digit-level accuracy (y-axis, hollow points) versus validation loss (x-axis) for a 6-digit addition task.

### v2

@@ F.1 Generation Process of the LM Training Data in Calcutec
     Reject the sampled $X\cup\{x^{\prime}\}$. $\triangleright$ We don’t want a premise that entails everything.
@@
   Sample a $g\in M^{\prime}$ such that the proof tree for $X^{\prime}\models g$ contains $s$ and its depth $d_{min}$. $\triangleright$ We use $d_{min}=4$ in our experiments.
@@
  $\leftarrow DD$.replace($s^{\prime}$, $r_{b}$)
@@
Algorithm 1 Pseudo code for the generation process of a Calcutec document used for training.
F.2 Perturbations in Calcutec
We apply two types of perturbations over the reasoning steps in Calcutec described in §7.1:
@@ F.2 Perturbations in Calcutec
Additionally, when flattening the proof trees with topological sort, we break the ties randomly. We also randomize the order of the symbols in the antecedents.
(a) document (tokens)
(b) paragraph (tokens)
(c) paragraph (steps)
(d) first $z$ (steps)
(e) document (tokens)
(f) paragraph (tokens)
(g) paragraph (steps)
(h) first $z$ (steps)
@@
F.3 Extra Setups
@@
Based on Assumption 6.1 and the formalism of NLP tasks in §3.2, input-label pairs of a downstream task corresponds to prefix-reference pairs in a paragraph. To examine whether the trained model can generalize well when the induction process for the label is different from the induction process for the pronoun in the training data, we generate a training dataset where all the pronouns are induced from the premise with a left-branching proof tree with a depth equal to 2 (Figure 4(a)), while the test data contains samples whose labels are induced from the input with balanced trees (Figure 4(b)).
@@
As described in §F.2, we apply some random perturbations on the proving process. We also experiment with the setup where we do not apply any perturbations.
@@
As described in §F.1, we rewrite the first step of the proof. We also experiment with the setup where we do not rewrite the first step.
@@ F.3.1 Results and Discussion
Figure 9 show that deeper models have better ICL performance. It aligns with the real-world observation that scaling helps model performance.
                                      $x_{1}$$x_{2}$$x_{4}$ $x_{3}$$r_{1}$ $x_{1}$$x_{3}$$x_{6}$ $x_{7}$ $\vdots\vdots$
@@ F.4 Hyper-parameters
(a) balanced, length = 44, $y_{1},y_{2}$
(b) branching, length = 44, $y_{1},y_{2}$
(c) branching, length = 33, $y_{1},y_{2}$
(d) balanced, length = 33, $y_{3},y_{4}$
(e) branching, length = 44, $y_{3},y_{4}$
(f) branching, length = 33, $y_{3},y_{4}$
Figure 5: In-context learning accuracy with Calcutec when using different verbalizers ($y_{1},y_{2}$ or $y_{3},y_{4}$) and input lengths (3 or 4). The dotted lines represent the performance of unseen combinations described in §7.1.2, while the different colors represent the number of formulas each class ($v_{+}$ or $v_{-}$) is associated to. The main lines represent the average accuracy of 5 tasks. We plot the performance of each task in lighter colors.
(a) balanced, length = 44, $r_{1},r_{2}$
(b) branching, length = 44, $r_{1},r_{2}$
(c) branching, length = 33, $r_{1},r_{2}$
(d) balanced, length = 44, $r_{3},r_{4}$
(e) branching, length = 44, $r_{3},r_{4}$
(f) branching, length = 33, $r_{3},r_{4}$
@@
(a) balanced, length = 44, $r_{1},r_{2}$
(b) branching, length = 44, $r_{1},r_{2}$
(c) branching, length = 33, $r_{1},r_{2}$
(d) balanced, length = 44, $r_{3},r_{4}$
(e) branching, length = 44, $r_{3},r_{4}$
(f) branching, length = 33, $r_{3},r_{4}$
@@
(a) balanced, $r_{1},r_{2}$, 25000 examples
(b) balanced, $r_{1},r_{2}$, 10000 examples
(c) balanced, $r_{1},r_{2}$, 5000 examples
(d) balanced, $r_{3},r_{4}$, 25000 examples
(e) balanced, $r_{3},r_{4}$, 10000 examples
(f) balanced, $r_{3},r_{4}$, 5000 examples
(g) branching, $r_{1},r_{2}$, 25000 examples
(h) branching, $r_{1},r_{2}$, 10000 examples
(i) branching, $r_{1},r_{2}$, 5000 examples
(j) branching, $r_{3},r_{4}$, 25000 examples
(k) branching, $r_{3},r_{4}$, 10000 examples
(l) branching, $r_{3},r_{4}$, 5000 examples
Figure 8: In-context learning accuracy with different sizes of
(a) balanced, length = 44, 3 layers
(b) balanced, length = 44, 4 layers
(c) balanced, length = 44, 5 layers
(d) branching, length = 44, 3 layers
(e) branching, length = 44, 4 layers
(f) branching, length = 44, 5 layers
@@
branching balance
$r_{1},r_{2}$ $r_{3},r_{4}$ $r_{1},r_{2}$ $r_{3},r_{4}$
#-shot 2 4 6 2 4 6 2 4 6 2 4 6
single 49.1 89.5 84.0 59.5 92.0 86.9 58.5 86.2 85.5 50.3 90.3 89.9
double 47.8 71.4 75.6 53.1 77.1 86.1 49.1 70.4 69.0 50.5 75.8 79.4
triple 46.7 65.7 70.7 50.6 73.4 79.4 46.0 60.2 61.4 49.8 67.0 70.4
Table 6: The CoT performance with 2, 4, or 6 examples.
Sample $N=\{n_{1},n_{2},n_{3},n_{4},n_{5}\}$ from either $\{0,1,\cdots,9\}$ or $\{6,7,\cdots,15\}$.
Sample four operators $O=\{o_{1},o_{2},o_{3},o_{4}\}$ from $\{+,\times\}$.
Initialize an empty list $S$ for the storing the steps.
while $\|N\|>1$ do
  Randomly sample an index $i$ of an operator from $O$. If $\times\in O$, then make sure that the sampled index corresponds to $\times$.
  Create a step $s\leftarrow[n_{i},n_{i+1},n_{i}o_{i}n_{i+1}\mod 16]$ and append $s$ to $S$.
  Remove $n_{i}$ and $n_{i+1}$ from $N$.
  Remove $o_{i}$ from $O$.
  Insert $n_{i}o_{i}n_{i+1}\mod 16$ into NN at position ii.
end while
return $S$.
Algorithm 2 Pseudo code for the generation process of a training example for the arithmetic task.
Given a list of steps $S=\langle s_{1},s_{2},\cdots,s_{4}\rangle$. (Each step is represented with a list, where the last number in the list is the right-hand side of an equation.)
Initialize n list $S^{\prime}=\langle s_{1}\rangle$ for the storing the randomly merged steps.
for $i\in\{2,3,4,5\}$ do
  Uniformly sample a number $r$ from $\{0,1\}$.
  if $r=0$ then
   Append $s_{i}$ to $S^{\prime}$.
  else
   if $S^{\prime}[-1][-1]\in s_{i}[:-1]$ then
     Remove $S^{\prime}[-1][-1]$ from $s_{i}$.
     $S^{\prime}[-1]\leftarrow\text{concatenate}(S^{\prime}[-1][:-1],s_{i})$
   end if
  end if
end for
return $S^{\prime}$.
Algorithm 3 Pseudo code for the generation process of a training example for the arithmetic task.
Appendix G Details of the Arithmetic Task
G.1 Training Data
When generating a training example, we first sample five numbers. These five numbers are either from $\{0,\cdots,9\}$ or $\{6,\cdots,15\}$. We then repeat the following process until there is only one number remaining in the list: We randomly pick two neighbor numbers from the list of five numbers and sum them and insert the result back into the list. This process results in a sequence of four steps, each of which is an equation that adds two numbers and produces one number. We outline the algorithm in Algorithm 2.
After we generate the four steps, we randomly merge steps into one step. We iterate through the steps. If the left-hand side current step contains the result of a previous step, we merge these two step into a single step at probability 0.5. For example, if the two steps are “1 + 2 = 3 . 3 + 4 = 7.” We merge these two steps as “1 + 2 + 4 = 7.” We outline the algorithm in Algorithm 3.
G.2 Hyper-parameters
We generate 80000 examples for training, 10000 for validation and 10000 for evaluation in the seen setup. For the unseen setup, we generate $6^{5}=7776$ examples. We train five Transformer models of the GPT-2-small architecture with five random seeds for 15 epochs. We use batch size 64 and the default training hyper parameters in HuggingFace transformers-4.46.3, i.e., with learning rate 5e-5, warm-up ratio 0, AdamW optimizer.
Appendix H Dataset License
  • SST-2: MIT
  • MR: unavailable
  • CR: unavailable
  • Subj: unavailable